# Speculating LLMs’ Chinese Training Data Pollution from Their Tokens

### v1

@@ Abstract
Tokens are basic elements in the datasets for LLM training. It is well-known that many tokens representing Chinese phrases in the vocabulary of GPT (4o/4o-mini/o1/o3/4.5/4.1/o4-mini) are indicating contents like pornography or online gambling. Based on this observation,
@@
(3) We study how to speculate training data pollution via PoC tokens’ appearances (token ID). Experiments on GPT and other 23 LLMs indicate that PoC tokens widely exist while GPT’s vocabulary behaves the worst: more than 23% long Chinese tokens (i.e., a token with more than two Chinese characters) are either porn or online gambling. We validate our speculation method on famous pre-training datasets like C4 and Pile. Then, considering GPT-4o, we speculate that the ratio of “波*野结衣”111a famous Japanese porn star’s name in Chinese and also a GPT-4o’s token with index 185,946. The full name is partially masked here for privacy concerns. related webpages in its training data is around 0.5%.
@@ 1 Introduction
Our key findings are as follows. By detecting 9 vocabularies of 23 LLMs, we find that PoC tokens widely exist. By estimation and verification, we find that Chinese corpus in open-sourced datasets like mC4 Xue et al. (2020) is polluted: 2-3% Chinese contents are polluted. In the end, by taking token “波*野结衣” and its three subsequence tokens (Figure 1) as an example, we speculate that related Chinese websites may take 0.5% of the whole Chinese pre-training dataset of GPT-4o  [^4] Code is open-sourced at: pollutedtokens.site.
@@ 5.1 PocTrace: trace Chinese data pollution
The argument above establishes theoretical upper and lower bounds. To confirm that these bounds are the supremum and infimum, we need to demonstrate that they are the smallest upper bound and the largest lower bound. We employ a proof by contradiction: we construct a naive training corpus “ab ab”. For the token “ab”, its frequency F(ab)=2F(``ab")=2 and its upper bound $\min_{\begin{subarray}{c}t_{k}\subseteq_{\text{sub}}``ab"\end{subarray}}F(t_{k})=\min\{F(``a"),F(``b")\}=\min\{2,2\}=2$. Suppose, for contradiction, that $\min_{\begin{subarray}{c}t_{k}\subseteq_{\text{sub}}``ab"\end{subarray}}F(t_{k})$ is not the smallest upper bound, this would imply the existence of another upper bound strictly smaller than $\min_{\begin{subarray}{c}t_{k}\subseteq_{\text{sub}}``ab"\end{subarray}}F(t_{k})=2$ yet still greater than or equal to F(ab)=2F(``ab")=2, leading to a contradiction; Similarly, for the token “a”, the frequency F(a)=2F(``a")=2 and its lower bound $\max_{\begin{subarray}{c}"a"\subseteq_{\text{sub}}t_{j}\end{subarray}}F(t_{j})=\max\{F(``ab")\}=2$. Assuming $\max_{\begin{subarray}{c}"a"\subseteq_{\text{sub}}t_{j}\end{subarray}}F(t_{j})$ is not the largest lower bound would imply the existence of another lower bound strictly greater than $\max_{\begin{subarray}{c}"a"\subseteq_{\text{sub}}t_{j}\end{subarray}}F(t_{j})=2$ yet still smaller than or equal to F(a)=2F("a")=2, again resulting in contradiction. Hence, these bounds indeed represent the supremum and infimum.
@@ 6 Conclusion
ACL Ethics Policy is respected in this work. This work studies pollutePd Chinese tokens within LLMs vocabulary and polluted content within Chinese train corpus. We investigate open-sourced LLMs vocabulary and train corpus whose terms, conditions, and copyright are respected. This paper may include offensive and upsetting content which need to be use with caution for future research.
@@ A.2 Labeling team
Based on our interdisciplinary research team with 6 experts owning PhD degrees of philosophy, sociology, Chinese linguistics, computer science, and artificial intelligence, we further build a labeling team by including 6 undergraduates from top-tier Chinese universities. So, there are 12 well-educated Chinese native speakers involved in the labeling process. This labeling team consists of 6 males and 6 females, aged between 19 and 40. We make this labeling team to avoid any bias due to education level, gender, or age. The labeling process for all long Chinese tokens (a token representing more than 2 tokens) takes more than 6 hours for all team members. Considering GPT-4o is a model built by OpenAI which is located in the USA, we use google.com as the default search engine to find the token-related web contents for determining labels.

### v2

@@ Abstract
Tokens are basic elements in the datasets for LLM training. It is well-known that many tokens representing Chinese phrases in the vocabulary of GPT (4o/4o-mini/o1/o3/4.5/4.1/o4-mini) [^1] Same for GPT-5 and -oss (August 26, 2025). are indicating contents like pornography or online gambling. Based on this observation,
@@
(3) We study how to speculate training data pollution via PoC tokens’ appearances (token ID). Experiments on GPT and other 23 LLMs indicate that PoC tokens widely exist while GPT’s vocabulary behaves the worst: more than 23% long Chinese tokens (i.e., a token with more than two Chinese characters) are either porn or online gambling. We validate our speculation method on famous pre-training datasets like C4 and Pile. Then, considering GPT-4o, we speculate the ratio of “波*野结衣” [^2] A Japanese porn star’s name in Chinese (partially masked for privacy concerns) and also a token with index 185,946. related webpages in its training data is around 0.5%.
@@ 1 Introduction
Our key findings are as follows. By detecting 9 vocabularies of 23 LLMs, we find that PoC tokens widely exist. By estimation and verification, we find that Chinese corpus in open-sourced datasets like mC4 Xue et al. (2020) is polluted: 2-3% Chinese contents are polluted. In the end, by taking token “波*野结衣” and its three subsequence tokens (Figure 1) as an example, we speculate that related Chinese websites may take 0.5% of the whole Chinese pre-training dataset of GPT-4o  [^5] Project website: https://pollutedtokens.github.io.
@@ 5.1 PocTrace: trace Chinese data pollution
The argument above establishes theoretical upper and lower bounds. To confirm that these bounds are the supremum and infimum, we need to demonstrate that they are the smallest upper bound and the largest lower bound. We employ a proof by contradiction: we construct a naive training corpus “ab ab”. For the token “ab”, its frequency F(ab)=2F(``ab")=2 and its upper bound $\min_{\begin{subarray}{c}t_{k}\subseteq_{\text{sub}}``ab"\end{subarray}}F(t_{k})=\min\{F(``a"),F(``b")\}=\min\{2,2\}=2$. Suppose, for contradiction, that $\min_{\begin{subarray}{c}t_{k}\subseteq_{\text{sub}}``ab"\end{subarray}}F(t_{k})$ is not the smallest upper bound, this would imply the existence of another upper bound strictly smaller than $\min_{\begin{subarray}{c}t_{k}\subseteq_{\text{sub}}``ab"\end{subarray}}F(t_{k})=2$ yet still greater than or equal to F(ab)=2F(``ab")=2, leading to a contradiction; Similarly, for the token “a”, the frequency F(a)=2F(``a")=2 and its lower bound $\max_{\begin{subarray}{c}``a"\subseteq_{\text{sub}}t_{j}\end{subarray}}F(t_{j})=\max\{F(``ab")\}=2$. Assuming $\max_{\begin{subarray}{c}``a"\subseteq_{\text{sub}}t_{j}\end{subarray}}F(t_{j})$ is not the largest lower bound would imply the existence of another lower bound strictly greater than $\max_{\begin{subarray}{c}``a"\subseteq_{\text{sub}}t_{j}\end{subarray}}F(t_{j})=2$ yet still smaller than or equal to F(a)=2F(``a")=2, again resulting in contradiction. Hence, these bounds indeed represent the supremum and infimum.
@@ 6 Conclusion
ACL Ethics Policy is respected in this work. This work studies polluted Chinese tokens within LLMs vocabulary and polluted content within Chinese train corpus. We investigate open-sourced LLMs vocabulary and train corpus whose terms, conditions, and copyright are respected. This paper may include offensive and upsetting content which need to be use with caution for future research.
@@
Acknowledgements
This work was supported by the Key Program of the National Natural Science Foundation of China under No. 62132011, the National Science Foundation for Distinguished Young Scholars under No. 62425201, Ant Group, and the Center of High Performance Computing, Tsinghua University.
@@ A.2 Labeling team
Based on our interdisciplinary research team with 6 experts owning PhD degrees of philosophy, sociology, Chinese linguistics, computer science, and artificial intelligence, we further build a labeling team by including 6 undergraduates from top-tier Chinese universities, including 12 well-educated Chinese native speakers, 6 males and 6 females, aged between 19 and 40. We make this labeling team to avoid any bias due to education level, gender, or age. The labeling process for all long Chinese tokens (a token representing more than 2 tokens) takes more than 6 hours for all team members. And they are paid six dollars per hour, which exceeds the minimum wage requirements. Considering GPT-4o is built by OpenAI in USA, we use google.com as the default search engine to find the token-related web contents for determining labels.